# Discrete Chi-Square Method can model and forecast complex time series, like El Nino data between 1870 and 2024

### v6

@@ Introduction
The modelling and forecasting of El Niño is considered notoriously difficult [55, 42, 6, 26, 35, 38, 53, 43]. We will show that our DCM can easily model and forecast the El Niño time series. Since the challenging analyses of all simulated extremely complex time series succeed (Sections 3.1-3.7), there are no rational reasons why our analysis of the simple El Niño time series would fail.
@@
Our El Niño data are from The Seasonal Mean Nino 4 HadISST1.1 (NOAA PSL) sample. These NOAA data were retrieved on January 29th, 2026. The twelve monthly sea surface temperature anomaly values give
@@
There were no values for the last four months of 2025. The full data contains $n=155$ yearly values between 1870 and 2024.
@@ Results
The DCM can model and forecast the El Niño time series. This result may be the “holy grail” of climatology.
@@
We detect significant signals only from the weighted data. Therefore, the trend and the forecasts are solved from the weighted data. The weighted DCM analysis results for all data are given in Table A1. For one, two, three, and four signals, the best trend alternative is the linear trend $K_{3}=1$. We use this $K_{3}=1$ trend in our DCM analysis of El Niño data. For the $\mathcal{M}$=3 model, the positive linear trend is $2M_{1}=0.56\pm 0.20$ $\mathrm{C}^{o}$ during $\Delta T=154$ years (Equation 5). This trend is approximately the same in all $\mathcal{M}$=1-4 models.
@@
We use the first half of all data $(n=78)$ to forecast the second half $(n=77)$. The results of weighted DCM analysis are given in Table A2. The periods of the three strongest signals are $P_{1}=5.580\pm 0.085$, $P_{2}=12.82\pm 0.40$ and $P_{3}=19.30\pm 0.83$ years (Model $\mathcal{M}$=3). The forecasts of the two and three signals models, open circles, are shown in Figures A1 and A2. For the second half of the data, the forecast $\chi^{2}$ values for the one, two, three and four signal models are 421, 336, 370 and 428, respectively. Hence, the two signal $\mathcal{M}$=2 model gives the best forecast.
@@
The DCM analysis results for all weighted data are given in Table A3. We detect the same periods that are already detected from the first half of all data. The three signal $\mathcal{M}$=3 model is the best. The El Niño signal periods are $P_{1}=5.662\pm 0.077$, $P_{2}=12.78\pm 0.12$ and $P_{3}=21.3\pm 1.5$ years. The DCM periodogram of the $\mathcal{M}$=3 model shows three clearly defined minima (Figure A3a). These three signals have the peak to peak amplitudes $A_{1}=0.40\pm 0.13$ $\mathrm{C}^{o}$, $A_{2}=0.47\pm 0.10$ $\mathrm{C}^{o}$ and $A_{3}=0.37\pm 0.12$ $\mathrm{C}^{o}$ (Figure A3b). The four signal $\mathcal{M}$=4 model is unstable $({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{UM}}})$. The three and four signal models are shown in Figures A4 and A5. For both alternatives, the predicted values after 2024, open circles, are given in Table A4. Both model alternatives give essentially the same forecast.
For the sake of consistency, we also give the results of non-weighted analysis of all data (Table A5). We detect the same signals from these non-weighted data, but the signal significances $Q_{\mathrm{F}}$ are lower. This confirms that the data error information of weighted data is crucial.
For all weighted data, the detection of the first signal is absolutely certain. The Fisher-test comparison between the $K_{3}=0$ pure noise model $g_{0,1,0,\chi^{2}}$ and the one signal model $g_{1,1,1,\chi^{2}}$ gives signal significance $Q_{\mathrm{F}}=3.9\times 10^{-13}$ (Table A6, $\mathcal{M}$=1). All other polynomial model alternatives, $1\leq K_{3}\leq 7$, are also rejected. The second and third signal significances are $Q_{\mathrm{F}}=2.7\times 10^{-7}$ and $Q_{\mathrm{F}}=0.00038$ (Table A3: $\mathcal{M}$=2-3).
@@
The DCM can model and forecast the El Niño phenomenon. The first half data forecast fits to the second half data (Figures A1b-d and A2b-d: open circles). For the two times larger all data sample, time will tell if our next decade forecast is also correct (Figures A4d and A5d: open circles). For example, we forecast an extreme El Niño between 2030 and 2032.
The solar forcing is the most probable explanation for these results. The three detected El Niño signal periods are close to the 22 years Hale cycle period [25], the 11 years solar cycle period [49, 59] and half the solar cycle period. A stochastic and non-stationary solar cycle [8, 57] can not cause these strict El Niño periodicities. Only very regular solar energy output changes can cause these El Niño signals because the noise added by the terrestrial climate response can not prevent our signal detection from the sea surface temperature data. These strictly periodic El Niño signals support the claim that the planets cause a stationary, multi-periodic and deterministic solar cycle(-s) [30].
Our mathematical model can detect regularities that the current physical models fail to detect, let alone forecast. Mathematics detects signals that should not be there according to Physics. The counterargument is that simple mathematics can not explain a complex physical phenomenon. We argue that the El Niño physical model is simple. The solar energy output determines El Niño. It is no surprise that the Sun warms the largest battery on the Earth, the Pacific Ocean. And if many cycles moderate the solar energy output, it is no surprise either that we have failed to discover the interference of these cycles, El Niño, in our yearly climate temperatures.
The current El Niño ”tool kit” is admittedly impressive [6]. Generations of hard work has given us those tools. Masses of scientists develop and use them. Yet, even the best currently available physical El Niño forecasts fail in 1.5 years [62]. Our mathematical DCM forecast can connect seven El Niño decades to the next seven El Niño decades (Figures A1 and A2). That is quite a coincidence if the chaos of physical models takes over in 1.5 years. The annual cost of El Niño damage is roughly one trillion dollars [9, 34]. An accurate El Niño prediction for just one year ahead would save the global economy trillions by mitigating long-term productivity losses [27, 61]. This should motivate other scientists to check our results immediately. [^4] All necessary file are in
@@
No supercomputers are needed because an ordinary PC can detect the first two signals in about one minute. The third signal detection takes a few minutes.
William of Ockham (1287–1347) stated that if two competing hypotheses make the same predictions, the simplest hypothesis is usually the best one (Occam’s razor). The connection between El Niño, the Sun and the Planets may seem like the holy grail. However, those three are just physical phenomena. The real scientific holy grail is the DCM.
@@ Acknowledgements
This work has made use of NASA’s Astrophysics Data System (ADS) services.
@@
            (c)             (d)
Figure A1: First half forecast (Table A2: $\mathcal{M}$=2): Forecasting data (closed black circles) and forecasted data (open black circles). Green and red dottted lines denote one and three sigma model $g(t)$ errors. Residuals (closed and open blue circles) are offset to level of -3. Units are x-axis: $t_{i}]$=y, y-axis: $y_{i}]=$ $\mathrm{C}^{o}$ and error bars = $1\times\sigma_{i}]=$ $\mathrm{C}^{o}$.
@@
            (c)             (d)
Figure A2: First half forecast (Table A2: $\mathcal{M}$=3). Otherwise as in Figure A1.
Figure A3: All data model $\mathcal{M}$=3 in Table A3. (a) Periodogram (Equation 21). Otherwise as in Figures 2ab. (b) Three $h_{1}(t)$, $h_{2}(t)$ and $h_{3}(t)$ signals as a function of time $t$ (left-hand panels) and phase $\phi$ (right-hand panels) [30, Equation 29].
@@
            (c)             (d)
Figure A4: All data model $\mathcal{M}$=3 in Table A3. Otherwise as in Figure A1.
@@
            (c)             (d)
Figure A5: All data model $\mathcal{M}$=4 in Table A3. Otherwise as in Figure A1.
Table A1: Trend in all weighted data. DCM analysis between $P_{\mathrm{min}}=3$ and $P_{\mathrm{max}}=100$ years. Otherwise as in Table 8
@@
Table A2: Periods in first half data. Otherwise as in Table A1.
@@
Table A3: Periods in all weighted data. Otherwise as in Table A1.
@@
Table A4: Forecasts (Figures A4 and A5: open circles).
Figure A4 Figure A5
@@
2025 0.09 0.24 0.35 0.28
2026 0.35 0.20 0.47 0.29
2027 0.41 0.21 0.31 0.28
2028 0.35 0.22 0.21 0.28
2029 0.40 0.19 0.42 0.34
2030 0.60 0.20 0.79 0.31
2031 0.79 0.20 0.94 0.21
2032 0.75 0.27 0.66 0.21
2033 0.45 0.34 0.18 0.23
2034 0.11 0.30 -0.10 0.31
2035 -0.03 0.21 -0.03 0.38
2036 0.06 0.22 0.16 0.34
Table A5: Periods in all non-weighted data. Otherwise as in Table A1.
@@
Table A6: Significance one signal model $\mathcal{M}$=1 in Table A3. Otherwise as in Table 9. 44footnotetext: Note: No test for $\mathcal{M}$=5  $(\eta_{1}=\eta_{2}=5)$.

### v7

@@ Introduction
The modelling and forecasting of El Niño is considered notoriously difficult [55, 42, 6, 26, 35, 38, 53, 43]. We will show that our DCM can model and forecast the “Big Wave” in El Niño time series. Since the challenging analyses of all simulated extremely complex time series succeed (Sections 3.1-3.7), there are no rational reasons why our analysis of the simple El Niño time series would fail.
@@
Our El Niño data are from The Seasonal Mean Nino 4 HadISST1.1 (NOAA PSL) sample. These NOAA1444https://psl.noaa.gov/data/timeseries/month/
data/nino4.long.anom.dat
data were retrieved on January 29th, 2026. The twelve monthly sea surface temperature (SST) anomaly values give the yearly values
@@
Four monthly values for the year 2025 were missing. Therefore, we analyse $n=155$ yearly values between 1870 and 2024.
@@ Results
The DCM can model and forecast the “Big Wave” in El Niño time series. This result may be the “holy grail” of climatology.
@@
We detect significant signals only from the weighted yearly data. Therefore, the trend and the forecasts are solved from the weighted data. The weighted DCM analysis results for all yearly data are given in Table S1. For one, two, three, and four signals, the best “Big Wave” alternative is the linear trend $K_{3}=1$. We use this $K_{3}=1$ trend in our DCM analysis of El Niño data. For the $\mathcal{M}$=3 model, the positive linear trend is $2M_{1}=0.56\pm 0.20$ $\mathrm{C}^{o}$ during $\Delta T=154$ years (Equation 5). This trend is approximately the same in all $\mathcal{M}$=1-4 models. This pattern supports climate change-induced global warming.
@@
We use the first half of all yearly data $(n=78)$ to forecast the second half $(n=77)$. The results of weighted DCM analysis are given in Table S2. The periods of the three strongest signals are $P_{1}=5.580\pm 0.085$, $P_{2}=12.82\pm 0.40$ and $P_{3}=19.30\pm 0.83$ years (Model $\mathcal{M}$=3). The forecasts of the two and three signals models, open circles, are shown in Figures S1 and S2. For the second half of the data, the forecast $\chi^{2}$ values for the one, two, three and four signal models are 421, 336, 370 and 428, respectively. Hence, the two signal $\mathcal{M}$=2 model gives the best “Big Wave” forecast.
@@
The DCM analysis results for all weighted yearly data are given in Table S3. We detect the same “Big Wave” periods that are already detected from the first half of all data. The three signal $\mathcal{M}$=3 model is the best. The El Niño  “Big Wave” signal periods are $P_{1}=5.662\pm 0.077$, $P_{2}=12.78\pm 0.12$ and $P_{3}=21.3\pm 1.5$ years. The DCM periodogram of the $\mathcal{M}$=3 model shows three clearly defined minima (Figure S3a). These three signals have the peak to peak amplitudes $A_{1}=0.40\pm 0.13$ $\mathrm{C}^{o}$, $A_{2}=0.47\pm 0.10$ $\mathrm{C}^{o}$ and $A_{3}=0.37\pm 0.12$ $\mathrm{C}^{o}$. The four signal $\mathcal{M}$=4 model is unstable $({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{UM}}})$. The three and four signal models are shown in Figures S3 and S4. For both alternatives, the predicted values after 2024, open circles, are given in Table S4. The $\mathcal{M}$=3 and $\mathcal{M}$=4 model alternatives give essentially the same “Big Wave” forecast between 2025 and 2036. We might as well have forecasted the “Big Wave” for the next century because our first half data forecasts for the second half data are quite accurate (Figures S1 and S2).
For the sake of consistency, we also give the “Big Wave” results of non-weighted analysis of all data (Table S5). We detect the same signals from these non-weighted data, but the signal significances $Q_{\mathrm{F}}$ are lower. This confirms that the data error information of weighted data is crucial.
For all weighted data, the detection of the first “Big Wave” signal is absolutely certain. The Fisher-test comparison between the $K_{3}=0$ pure noise model $g_{0,1,0,\chi^{2}}$ and the one signal model $g_{1,1,1,\chi^{2}}$ gives signal significance $Q_{\mathrm{F}}=3.9\times 10^{-13}$ (Table S6, $\mathcal{M}$=1). All other polynomial model alternatives, $1\leq K_{3}\leq 7$, are also rejected. The second and third “Big Wave” signal significances are $Q_{\mathrm{F}}=2.7\times 10^{-7}$ and $Q_{\mathrm{F}}=0.00038$ (Table S3: $\mathcal{M}$=2-3).
@@
The DCM can model and forecast the “Big Wave” of El Niño. The first half data forecast fits to the second half data (Figures S1b-d and S2b-d: open circles). For the two times larger all data sample, time will tell if our next decade “Big Wave” forecast is also correct (Figures S3d and S4d: open circles). For example, we forecast an extreme El Niño between 2030 and 2032.
The solar forcing is the most probable explanation for this El Niño  “Big Wave” result. The three detected El Niño signal periods are close to the 22 years Hale cycle period [25], the 11 years solar cycle period [49, 59] and half the solar cycle period. A stochastic and non-stationary solar cycle [8, 57] can not cause these strict “Big Wave” periodicities. Only very regular solar energy output changes can cause these signals because the noise added by the terrestrial climate response can not prevent our signal detection from the sea surface temperature data. These strictly periodic El Niño  “Big Wave” signals support the claim that the planets cause a stationary, multi-periodic and deterministic solar cycle(-s) [30].
Our mathematical model can detect regularities that the current physical models fail to detect, let alone forecast. Mathematics detects signals that should not be there according to Physics. The counterargument is that simple mathematics can not explain a complex physical phenomenon. We argue that the El Niño physical model for the “Big Wave” is simple. The solar energy output changes could determine the “Big Wave”. It is no surprise that the Sun warms the largest battery on the Earth, the Pacific Ocean. And if many cycles moderate the solar energy output, it is no surprise either that we have failed to discover the interference of these cycles, the “Big Wave”, in our yearly climate temperatures. The immense Pacific Ocean surface acts like a giant thermometer that measures solar energy output in real time.
The current El Niño ”tool kit” is admittedly impressive [6]. Generations of hard work has given us those tools. Masses of scientists develop and use them. Yet, even the best currently available physical El Niño forecasts fail in 1.5 years [62]. Our mathematical DCM forecast can connect seven “Big Wave” decades to the next seven “Big Wave” decades (Figures S1 and S2). That is quite a coincidence if the chaos of physical models takes over in 1.5 years. The annual cost of El Niño damage is roughly one trillion dollars [9, 34]. An accurate El Niño prediction for just one year ahead would save the global economy trillions by mitigating long-term productivity losses [27, 61]. This should motivate other scientists to check our results immediately. [^5] All necessary file are in
@@
No supercomputers are needed because an ordinary PC can detect the first two “Big Wave” signals in about one minute. The third signal detection takes a few minutes.
William of Ockham (1287–1347) stated that if two competing hypotheses make the same predictions, the simplest hypothesis is usually the best one (Occam’s razor). The simple connection between El Niño  “Big Wave”, the Sun and the Planets may seem like the holy grail. However, those three are just physical phenomena. The real scientific holy grail is the DCM.
Note added to the proof
We arrive at a "Mexican standoff" because only future data can confirm that our “Big Wave” forecasts are correct. Fortunately, one minor detail saves us one year of waiting.
On January 29th, 2026, we computed the yearly values from the NOAA1 data. Four monthly values for the year 2025 were missing and our analysed data ended in the year 2024. Today February 25th 2026, NOAA1 still has these same four missing monthly values for the year 2025 (“-99.99”). To get these missing values, we are advised to use the Climate Prediction Center (CPC) data in NOAA2.666https://www.cpc-ncep.noaa.gov/products/
analysis-monitoring/ensostuff/ON-v5.php
All twelve monthly values for the year 2025 are given in Table S7 (column 2). These values give $y_{i}\pm 1\sigma_{i}=-0.26\pm 0.21$ $\mathrm{C}^{o}$. A red circle now highlights this observed new 2025 year value which is added into Figures S3e and Figures S4e. Our “Big Wave” forecast for the year 2025 succeeds. The year 2025 forecast for the best $\mathcal{M}$=3 model for all data (Table S3) is more accurate than the forecast for the unstable $\mathcal{M}$=4 model. Our forecast for this 2025 observation can, of course, be dismissed just as one lucky guess. Who wants to take the risk that our severe “Big Wave” forecasts for the years 2030 and 2031 are also “lucky”?
We also analyse the monthly SST values (Tables S8-S19, Figures S5-S16). The red dots in our figures highlight the available new monthly values between January 2025 and February 2026. Most of those red dots hit their mark. We consistently detect the “new” 3.6 and 4.6 year periods from these monthly values. The monthly mean computation eliminates the detection of these periods from yearly data. We conclude that the chaotic $\pm 1$$\mathrm{C}^{o}$ weather changes can prevent accurate DCM monthly forecasts. However, we can compute the “Big Wave” monthly values (Table S20). We admit that this does not fully solve the El Niño riddle but it may help. If no internal mechanism can cause the strictly multi-periodic “Big Wave” in the Pacific Ocean then the Sun is the only possible external cause. The planets may indeed cause regular solar cycles [30].
@@ Acknowledgements
This work has made use of NASA’s Astrophysics Data System (ADS) services. We thank Dr. Jouni Räisänen for his comments of the manuscript.
@@
            (d)             (e)
Figure S1: First half forecast (Table S2: $\mathcal{M}$=2): (a) Periodogram (Equation 21). (b-e) Forecasting data (closed black circles) and forecasted data (open black circles). Green and red dottted lines denote one and three sigma model $g(t)$ errors. Residuals (closed and open blue circles) are offset to level of -3. Units are x-axis: $t_{i}]$=y, y-axis: $y_{i}]=$ $\mathrm{C}^{o}$ and error bars = $1\times\sigma_{i}]=$ $\mathrm{C}^{o}$.
@@
            (d)             (e)
Figure S2: First half forecast (Table S2: $\mathcal{M}$=3). Otherwise as in Figure S1.
@@
            (d)             (e)
Figure S3: “Big Wave” for all yearly mean data model $\mathcal{M}$=3 in Table S3. Red circle is added observed 2025 year value. Otherwise as in Figure S1.
@@
            (d)             (e)
Figure S4: “Big Wave” for all yearly mean data model $\mathcal{M}$=4 in Table S3. Otherwise as in Figure S2.
Table S1: Trend in all yearly mean weighted data. DCM analysis between $P_{\mathrm{min}}=3$ and $P_{\mathrm{max}}=100$ years. We use $\gamma=0.001$ (Equation 48) for yearly data. Otherwise as in Table 8
@@
Table S2: Periods in first half yaerly mean data. Otherwise as in Table S1.
@@
Table S3: Periods in all yearly mean weighted data. Otherwise as in Table S1.
@@
Table S4: “Big Wave”yearly forecasts (Figures S3d and S4d: open circles).
Figure S3 Figure S4
@@
2025 0.3 0.2 0.2 0.2
2026 -0.2 0.2 0.2 0.2
2027 -0.4 0.3 -0.0 0.2
2028 0.2 0.3 -0.1 0.2
2029 0.5 0.3 0.2 0.2
2030 0.3 0.3 0.4 0.2
2031 0.1 0.2 0.4 0.2
2032 0.2 0.3 0.2 0.2
2033 0.2 0.3 0.1 0.3
2034 0.0 0.2 0.0 0.2
2035 0.1 0.2 -0.1 0.3
2036 -0.0 0.3 -0.2 0.3
Table S5: Periods in all non-weighted yearly mean data. Otherwise as in Table S1.
@@
Table S6: Significance one signal model $\mathcal{M}$=1 in Table S3. Otherwise as in Table 9. 66footnotetext: Note: No test for $\mathcal{M}$=5  $(\eta_{1}=\eta_{2}=5)$.